Offshore operations

Expedition 313 carried three types of coring tools: a hydraulic piston corer (HPC; equivalent to the advanced piston corer [APC]), an extended nose corer (EXN; equivalent to the extended core barrel [XCB]), and a standard rotary corer (ALN; equivalent to the rotary core barrel [RCB]). See Table T1 for a summary of operations.

Mobilization of the L/B Kayd

Mobilization of the L/B Kayd began with shipping of the ESO laboratory, database, and office containers from ESO partner institutes in Europe. Once all the equipment and containers had cleared customs, they were delivered to the U.S. Coast Guard Station in Atlantic City, New Jersey (USA).

On 24 April 2009, the L/B Kayd arrived ahead of schedule at the Coast Guard Station in Atlantic City. Mobilization was started immediately by the drilling contractor, Drilling, Observation and Sampling of the Earth's Continental Crust (DOSECC), who loaded the drilling and ESO containers onto the platform. ESO staff from the British Geological Survey (BGS) arrived at the L/B Kayd on 28 April to prepare the containers for operations, including installing a shipboard computer network. Because of the nature of transiting in a lift boat, with water spilling over the deck, power could not be fully installed in the containers until the first hole was reached.

Mobilization continued for 6 days. The wireline logging equipment was transferred to the L/B Kayd out of Miller's Launch, Staten Island, New York (USA), on 1 May with the first supply boat and crew change.

Transit to Hole M0027A

At 1030 h on 30 April 2009, the L/B Kayd and Expedition 313 set sail from Atlantic City and headed for the first coring site (Hole M0027A). A brief stop was made on the way to Hole M0027A to modify the flooring of the cantilevered drilling platform, which was being forced up by waves breaking on the L/B Kayd's bow. At 2345 h, the L/B Kayd arrived in Hole M0027A and prepared to position above the first hole.

Hole M0027A (proposed Site MAT-1)

At 0003 h on 1 May 2009, the L/B Kayd was positioned above Hole M0027A and the legs were lowered to tag the seabed. Once the seabed was tagged, the preload procedure began by gradually increasing the load on the seabed. The preload procedure was interrupted to reposition the L/B Kayd so that the communication satellites were not eclipsed by the legs. The preload and settlement procedure continued until 1053 h, when the L/B Kayd was jacked-up to 30 ft above the water. Normal access to the working deck was granted, and all teams prepared to begin coring. At 1215 h, the supply vessel Rana Miller arrived at the platform and delivered equipment and personnel, which included the wireline logging tools and the remaining scientific staff. At 1640 h, the drill rig was started, the mast was raised, and the conductor pipe was run to the seabed, tagging it at 2134 h. The core barrel was lowered to just above the seabed and the first HPC core was fired at 2350 h.

The first core arrived on deck at 0010 h on 2 May. Coring continued using the HPC, and the hole was advanced by recovery. Sixteen HPC attempts resulted in two nonfires and a penetration of 28 m in just under 20 h. Three core runs showed signs of caving, with the amount of caving fill between 0.33 and 0.72 m. Once the base of the sand layer had been established, coring was stopped. At 2035 h, the drill string was tripped and the conductor pipe extended into the seabed to act as casing.

At 0130 h on 3 May, the conductor pipe was set at 17.6 m drilling depth below seafloor (DSF). The hole was conditioned, and the PQ drill string was run to the base of the hole where HPC coring commenced. After two HPC attempts, the coring tool was switched to the EXN because of a hard lithology preventing HPC recovery. However, recovery was still poor. The coring method was switched back to the HPC, upon which 6 m of cavings was encountered. In an attempt to stabilize the hole, the mud was changed and the hole was reamed with a full-face bit in a standard rotary coring barrel (ALN). Hole stability problems continued and the full-face bit was found to have no cutting surface left when retrieved at 1830 h. The string was tripped back to the surface to inspect the outer bit; however, no damage was observed. The string was rerun to the base of the hole, and coring continued with the HPC, which when retrieved contained a few pebbles and had a dented nose. The ALN was run to clear the hole shortly before midnight.

During the rerun of the ALN corer, the bit blocked again. A second ALN corer was run at 0100 h on 4 May, which on the second attempt advanced the full 3 m but recovered no core. At 0353 h, HPC coring advanced the hole but ended when the drill string became stuck in the hole. After the drill string was freed, it was recovered to the deck and checked for damage. The string was lowered back into the hole to within 9 m of the base, where it was stopped by infill material. It was also suspected that pieces of the previously broken full-face bit had been encountered, and the hole was reamed out with the ALN corer to save the PQ string bit. Coring recommenced using the EXN tool. At 1620 h, it was decided that the depth of casing in the hole should be increased. After pulling the PQ string, the casing could not be moved in the hole, so the hole was reentered using the PQ string and the hole advanced using the EXN tool. The EXN was used in favor of the HPC, as the HPC seemed to encourage hole collapse.

By 0400 h on 5 May, Hole M0027A was good, clean, and free-running with circulation at the base. EXN coring continued throughout the day with mixed recovery, although infill was encountered with most new core runs. The core barrel handling procedure was modified such that two EXN core barrels were always operational, which reduced the time the hole was left vulnerable to cave-ins. This, coupled with using the EXN, dramatically improved the coring. From 2000 h an increasingly hard lithology caused core runs to become blocked, and a liner was crushed at the shoe.

At 2355 h on 5 May, operations were halted because of a thunderstorm.

Coring restarted at 0255 h on 6 May using the ALN, but problems with crushed liners blocking the bit persisted throughout the first half of the day. At 1500 h, the mud mix was altered slightly, and the ALN corer switched for the HPC tool, which improved recovery slightly. At 1830 h a meeting was convened to discuss how to best advance the hole, as progress to date had been slow. All agreed to HPC spot core in maximum 60 ft increments (6 x 10 ft pipe lengths, ~18 m), or sooner if a variation in drilling parameters was encountered, suggesting a change in lithology. This was done to ~180 m DSF, where continuous coring resumed. The first open-hole interval commenced at 2140 h and continued past midnight.

Rapid progress was made at the beginning of 7 May using the strategy of HPC spot coring every 60 ft or less. Seven HPC attempts over 64 m were made in 20 h, with a slight setback occurring when the HPC tool became stuck in the bottom-hole assembly (BHA) and the wireline wire broke while trying to pull the tool free. The result of this was that the PQ string had to be pulled out of the hole, which was a wet pipe trip because of the HPC sealing the BHA. The wireline was not fished, as there was no way to circulate the mud during fishing, which would have increased the risk of hole collapse.

The wireline wire, which had broken ~80 ft above the BHA, was replaced early in the morning of 8 May, as was a worn bit on the BHA. The drill string was tripped back into hole with the noncoring bit inserted. The string reached the base of the hole 20 m higher than expected because of infill, which was subsequently drilled out. The noncoring barrel got stuck 3 m above the base of the hole when it was unlatched from the BHA. This necessitated another drill string trip. After 14 h, HPC coring resumed at 1325 h. The liner collapsed in several runs because of very stiff clay, which limited core recovery and caused several core barrels to become stuck in the BHA.

Problems with the HPC tool getting stuck in the BHA continued for the first 3 h of 9 May. After partly tripping the drill string, the HPC became free and was returned to the deck. From 0330 h, EXN and HPC coring were alternately conducted, with the EXN tool eventually chosen in favor of the HPC tool. Seven runs with the EXN tool collected very good core before chattering on the drill string, and no penetration indicated a change back to sand.

To tackle this lithology, the ALN corer was prepared, and early on 10 May coring resumed with the ALN in what appeared to be a gravel layer. There was no recovery in this layer. A mud backflow occurred, and large bubbles were observed rising up the drill string. No H2S or abnormal smell was recorded by the DOSSSEC or ESO instruments, and it was assumed that the gravel layer was hosting a freshwater flow. The next core did have an abnormal smell, and tests using the ESO gas analyzer measured 7.7 ppm H2S and no flammables. A later core tested positive (1%) for flammables near a thin, very dark layer. Excellent coring conditions prevailed into Monday, 11 May, through alternating hard clays and hard fine sands.

Progress slowed slightly when the lithology became more sandy, until the hole collapsed and the drill string became stuck at 1150 h. Upon freeing the string, 9 m of infill blocked the hole and an increase in rotational torque was observed. To reduce this pressure on the string, thought to be from the clay formation above, the string was pulled back a total of 26 m to ream the hole. Progress continued through very loose, clean sand for the rest of the day, with surprisingly excellent core recovery. On 13 May, very good progress continued using the ALN corer, reaching 451.06 m DSF at midnight.

By 0430 h on 14 May, it became apparent that the BHA stabilizer rings were worn and needed replacement. The string was tripped in order to replace both rings. The trip was completed by 0800 h, the rings replaced, and the pipe run back into the hole. The bottom of the hole was tagged at 335 m DSF because of a bridge that had formed, at which depth open holing commenced to clear it. A zone of high-pressure water was encountered, which continued until 345 m DSF. Below this, the backpressure decreased to normal and the string broke through the blockage. Infill was encountered after another 8–10 m of air gap beneath the bridge. Approximately 100 m of fill was drilled out to reach the bottom of the hole. Coring recommenced using the ALN corer at 1910 h.

Progress slowed on 15 May, although core recovery was generally good with near 100% for many of the runs. However, some sections of some cores were undersized. The ALN corer was replaced by the EXN corer for one run to see if it gave better results. The lithology proved to be unsuitable for EXN coring with little penetration, high backpressure, and torque, so the ALN corer was used for the rest of the day. In the evening, the core barrel got stuck in the BHA at a depth of 509 m DSF. It was eventually freed, but after this event core recovery was poor and there were ongoing problems with core barrel latching and retrieval. By the end of the day the depth of the hole was 515 m DSF.

Steady progress was made on 16 May with the ALN corer. Core slippage occurred on some core runs, but the core was usually retrieved on the next run. Good progress was made on 17 May, with most cores having 100% recovery. Repairs to the wireline and bit refurbishment caused some delays in the afternoon and early evening.

On 18 May, 11 core runs were made with >30 m penetration using the ALN corer. Recovery was more variable than in recent days (85%). In some cases this was due to core slippage, although slipped material was often recovered in the next run. The rate of penetration slowed because of a combination of the formation becoming harder and the round trip of the core barrel taking longer with increasing depth in the hole. There were reports of a "petroleum smell" from the drillers in the early morning. However, there was no reading on the gas monitors. A slight increase in torque was noticed in the evening.

Shortly before midnight on 18 May, the core barrel was recovered to deck with no core. It was suspected that the core barrel had not latched into the BHA. Another barrel was deployed and became jammed in the upper section of the drill string. This was discovered when the overshot latched onto the barrel at approximately the waterline. After removing 11 double stands of pipe, the overshot with the latch head was also unable to pass through, although the overshot alone could. The core barrel was deployed with the latching indicator ball installed and successfully latched into the BHA. After adding pipe (0410 h, 19 May) to return the string to the base of the hole, increased torque, backlash during rotation, and backpressure on the mud gauge was noted. When the string was broken to add further lengths, frothing and overflow of mud occurred. On lowering the overshot, slack on the wireline indicated it was being temporarily stopped in the drill string. The overshot latched on as normal, but on recovery the assembly became jammed at ~25 m below deck level. After two attempts to release the core barrel, it was recovered to deck. Scratches and spiral polishing were noted on the outside of the barrel.

At 0600 h on 19 May, the coring operation in Hole M0027A was stopped and preparations made to start the logging program. Because of the instability in the upper part of the hole, the decision was made to log the open hole in three sections. Through-pipe total gamma ray was acquired along the full length of the hole. Open-hole resistivity and magnetic susceptibility were acquired for the bottom section only, from 430 to 634 m DSF. Logging continued, running sonic velocity, acoustic imagery, and spectral gamma ray for the lower interval. The pipe was then pulled to 195 m DSF, beginning at 1455 h on 20 May, and resistivity, magnetic susceptibility, and spectral gamma ray were acquired between 195 and 347 m DSF. A bridge that had formed at 347 m DSF prevented logging of the section between 347 and 430 m DSF.

At dawn on 21 May, preparations were made to begin a vertical seismic profile (VSP), beginning with a marine mammal watch at 0513 h. Air gun firing began at 0548 h, and the VSP tool was inserted into the pipe at 0600 h. VSP work continued throughout the day until about 1900 h. Open-hole VSP data were acquired from 323 to 195 m DSF with through-pipe VSP data from 195 m DSF to seafloor. Marine mammal observations ceased at 1930 h.

Preparations were made to pull the pipe up to the top of the next open-hole logging section. However, a fault developed with the drilling rig motor before pipe pulling commenced. This required a spare part which was not on the platform. Efforts to make a temporary repair on board to keep the system operational for the remainder of the logging program were attempted but were unsuccessful. Late in the evening, operations ceased while waiting for the part (and additional spare) to be delivered by supply boat on 22 May.

The rig was operational by 0830 h on 22 May. Two double stands were tripped with the aim of opening the hole to 97 m DSF. However, the pipe became stuck, and despite applying five times the usual torque, the pipe would not rotate in the hole and could not be pulled further. The decision was taken to abandon the hole and move to Hole M0028A to begin coring operations. The pipe was cut just above the core barrel and tripped. The casing was retrieved and the deck, drilling rig and containers secured for the rig move.

Transit to Hole M0028A

Preparations for transit to Hole M0028A commenced at 2100 h on 22 May 2009. The jack-down procedure began at 2315 h, and the L/B Kayd moved off Hole M0027A at 2335 h, arriving at Hole M0028A at 0100 h on 23 May. The legs were lowered to the seafloor by 0120 h, when the preloading procedure began. The drilling floor was opened up to operations personnel at 1100 h to begin set-up of generators, powering up the containers, and preparation of the drilling floor.

Hole M0028A (proposed Site MAT-2)

The casing plan for Hole M0028A was to run the casing as deep in the hole as possible. The casing operation began after midday and continued overnight. Progress was very slow because of the ground conditions. Just before 0600 h on 24 May 2009, the casing twisted off at the crossover sub 1–2 m above the seabed. Remotely operated vehicle (ROV) inspection confirmed that casing was protruding from the seabed and lengths of casing pipe were lying on the seabed, which would require a salvage operation to remove.

The decision was taken to try and restart the hole at the same location. The new casing string was prepared and run into the seabed and, rather than case to great depth, it was decided to open-hole using the PQ string to the target depth of 215 m DSF and then continuously core the middle and lower sections of the hole to 750 m DSF, as this was the primary objective. By the end of 24 May, 80 m had been drilled with the noncoring inner barrel inserted.

At 0030 h on 25 May, recovery of the noncoring barrel (to flush the pipe) began. However, the drill pipe got stuck at a depth of 102 m DSF. The jam was cleared by 0830 h, and open-hole drilling resumed, reaching a depth of about 157 m DSF at 1450 h. Drilling stopped at this point to lower the casing further into the seabed using the L/B Kayd and then restarted. Just before 1800 h a clay formation was encountered at a depth of 178 m DSF. The core barrel was inserted, but the drill pipe was sticking, so the base of the borehole was reamed out. This proved unsuccessful. Sticking points had been encountered higher up in the hole (probably swelling clays), so a much more substantial reaming operation began in order to condition the hole before the start of continuous coring operations. The pipe was tripped back to 98 m DSF, and a mud mix that would help inhibit the swelling clays was pumped into the hole.

The reaming operation continued overnight and for most of the following morning. The hole was open-hole drilled to 218 m DSF by 1230 h on 26 May, and coring began using the ALN drill string. Backpressure was encountered during coring on the first and second runs. After that, backpressure diminished. A zone of high backpressure was encountered in Hole M0027A at a similar depth interval. Coring and reaming continued until midnight, with core recovery 100% or greater on all runs, reaching a depth of 246 m DSF. Overnight, core recovery became more variable as sandy formations were encountered and there were problems with the drill pipe sticking, which required more hole conditioning. At 0820 h on 27 May at a depth of 264 m DSF, the core barrel did not latch in when lowered. It could not be retrieved, and there was no flush. The only solution was to trip the pipe and recover the barrel. The pipe was tripped by 1330 h. The bottom of the BHA was blocked with sand. As the drill pipe had been tripped, this was an opportunity to try to deepen the casing beneath the seabed to improve its stability. The casing operation continued until midnight.

On 28 May, casing continued to be run, interrupted for a few hours by mechanical problems with the drilling rig. An ROV inspection at 0900 showed that the casing had entered the seabed close to the previously abandoned casing and clear of any casing pipe debris. The operation was completed at 1945 h, by which time the foot of the casing was at 22.39 m DSF and occupying the hole from which the PQ pipe and casing had been previously tripped. The operation to lower the PQ drill pipe back to 264 m DSF and to clean and ream the hole continued overnight and was completed by 1945 h on 29 May.

The first cores were very sandy with poor recovery. There were also problems getting the core barrel to latch in because of sand entering the BHA. This required removal of some pipes to raise the bit from the bottom of the hole and pumping mud to clear the BHA. Once flushed clear, the core barrel was lowered and latched in, and then the drill pipe was lowered to the base of the hole to begin coring again. This operation was completed by 2310 h. Backpressure due to artesian water was also noted in this zone.

By midnight on 29 May, the base of the hole had advanced to 270 m DSF. Coring continued throughout 30 May with the ALN core barrel, and by midnight the hole had advanced to 317.62 m DSF. There were some problems with recovery and blocked core barrels due mainly to alternating beds of lithified sandstones and unlithified silty sands and clays. Zones of high backpressure were also encountered at 274 and 305 m DSF, but there was no evidence for gas being the cause.

From about 0800 h on 31 May, recovery dropped significantly as a coarse, sandy formation was encountered. By 1200 h the bottom of the hole had advanced to 339 m DSF. At this time the drill pipe (probably the BHA) got stuck in the hole following a cave-in. After some effort, the drill pipe was freed, but it was apparent that there was a problem with the bit on the BHA. There was little further penetration, and on retrieval, the inner barrel showed signs of damage and scoring. The drill pipe was tripped, and the BHA was on deck at 1740 h. The bit and reaming shell were missing and were probably sheared off while freeing the drill pipe. Rather than attempt to restart drilling immediately with a new BHA, it was decided to fish for the missing parts, as they could impede further progress in the hole. The Bowen spear was made ready, and the fishing operation began at 2020 h.

The fishing operation proceeded very slowly throughout the night. At 0600 h on 1 June, the tool was at a depth of 198 m DSF and was encountering bridges that were difficult to wash away. Around 1200 h the tool encountered a zone of swelling clays that was known to extend to 240 m DSF. Progress was halted by the clays, torque on the string was very high, and the drill string was not advancing. At 1230 h the decision was taken to abandon the fishing operation, as there was little chance of further progress downhole and there was the danger that the tool would also be lost. The tool was tripped and was back on deck at 1430 h. A new BHA was prepared with a diamond impregnated drill bit, capable of drilling through or pushing aside the missing bit and reaming shell. By midnight on 1 June, washing and reaming down had advanced the drill string to 143 m DSF, and by midnight on 2 June it was at 326.77 m DSF.

Coring started at 0215 on 3 June and continued until midnight. The cored material was very variable for several core runs, from solid sandstone to running sand, with thin clay layers. The loose sand caused caving and slowed the progress for ~15 m, with several cores incorporating infill.

On 4 June, coring from 360 m DSF became easier, although the core recovery was variable. Small cavings occasionally slowed progress through alternating soft and firm formations. From 1750 h onward, sandy formations forced shorter core runs and caving sands required additional reaming.

At ~0100 h on 5 June, a bridge formed above the BHA that caused the drill string to become stuck. Several hours were spent trying to pull the BHA past the bridge and circulate mud. This was unsuccessful, and the decision was made at 0900 h to continue coring using the HQ drill string through the PQ drill string, effectively using the PQ string as casing set at 404 m DSF. Before the HQ string was run into the hole, a through-pipe gamma ray wireline log was taken from 390 m DSF to the surface, which was completed by 1530 h. A total of 10 m of infill material in the PQ pipe had to be reamed out before HQ coring could begin.

At 0330 h on 6 June, coring using the HQ Tight Tolerance (HQTT) core barrel commenced. After four core runs, a hole collapse due to fine caving sands risked the HQ string becoming stuck, and so the HQ string was pulled back into the PQ casing (0830–1320 h). Attempts were then made to ream back down, but the hole pressurized and four additional pipes had to be removed. A second attempt at reaming back down was successful, and coring started again at 1620 h. However, on the fifth HQTT core run, the drill string became stuck again for an hour. Amendments were subsequently made to the mud mixture.

The first few hours of 7 June were spent circulating and advancing the HQ string to the base of the PQ casing. While advancing the HQ string, the driller noted little penetration and thought there was an obstruction. The HQ string was pulled back to the surface and the BHA checked. The bit was still attached but severely worn and was replaced with a similar but harder matrix impregnated bit. Upon running the HQ string down the hole, 67 m of sand was encountered inside the PQ string. The remainder of the day and the beginning of Monday, 8 June, was spent reaming out this material. A core of infill was taken from material believed to represent ~403.79–411.86 m DSF.

At 0745 h on 8 June, normal coring commenced with nine good cores collected. The improvement in coring progress continued into Tuesday, 9 June. Coring was suspended from 1055 to 1225 h because of an electrical storm.

Coring continued to progress well until 1200 h on 10 June, with eight core runs and 18.5 m penetration, reaching a depth of 533.85 m DSF. However, a decreasing penetration rate caused concern, with the view that the inside cutting diameter of the core bit had worn out and thus the bit was cutting an oversized core. It was decided to pull the HQ string to check the bit. This decision coincided with the detection of H2S in the last core run returned to the deck (~1100 h). Precautions were taken, including venting the core prior to curation, and the borehole was monitored. The HQ string was pulled and the bit inspected. Although the bit showed considerable wear, it still did not account for the poor penetration. At 1725 h, the HQ string was run back into the hole with a polycrystalline diamond (PCD) bit attached. The hole appeared to be stable, and ~1.2 m of cavings was cleared. After reaming, the core barrel was retrieved for an H2S test (to 2400 h).

From 0035 to 0515 h on 11 June, there was a smell of H2S on the drill floor, although nothing registered on the gas monitors. The first core of the day had a smell of H2S and registered 4.3 ppm on the gas monitor, which soon dissipated. Drilling continued with regular checks for H2S returning with the mud. The second and third cores also had a smell of H2S but did not register on the gas monitor. Several more core runs were made, resulting in near 100% recovery, but failed to penetrate the full length of the 3 m core run because of a harder lithology. As the lithology changed to a more loose material, full core runs were recovered.

An electrical storm briefly halted operations from 0255 to 0330 h on 12 June. Penetration rates slowed in comparison to previous days, but with no discernible change in the lithology.

Early on 13 June, the drill string was pulled, and it was found that the outer cutting edge of the bit had worn away. At 0555 h the HQ string was run back down with a new bit and stabilizer ring. At 326 m DSF, very high mud backpressure indicated a blocked bit. It was not possible to latch the core barrel with the overshot. A second HQ trip was made, and sand was found behind the core barrel blocking its release. On running the HQ string back in, it was found that the PQ string had filled up with sand to 250 m DSF, and flushing was required. At ~439 m DSF, the drill string easily dropped down into the hole.

The HQ string reached the base of the hole at 0530 h on 14 June, and coring recommenced with the HQTT barrel. Core recovery varied throughout the day but improved slightly on 15 June when coring entered stiff, swelling clays.

At 0225 h on 16 June, the end of coring in Hole M0028A was declared. At 0240 h, the final core was curated. The last 10 double rods (61 m) were lifted to ream out the bottom of the hole in readiness for logging. The wireline and VSP logging winches were set up on either side of the drill rig. At 1140 h, through-pipe spectral gamma ray logging began and was completed by 1650 h. A marine mammal watch began at 1730 h prior to the through-pipe VSP logging commencing at 1800 h.

The VSP tool was returned to the deck by 0030 h on 17 June. At 0130 h, mud circulation began and the HQ string pulled to a clay layer ~25 m below the PQ string, keeping it below the running sand that caused earlier coring and reaming problems. Between 0450 and 2100 h, wireline logging and VSP of the lower open-hole interval from 674 to 451 m DSF was conducted with no problems. The air guns were left in the water, test firing until the next VSP run (through–PQ pipe) which started at ~0200 h on 18 June. The marine mammal observation continued throughout this period.

At midnight on 18 June, the HQ string was pulled slowly out of Hole M0028A. The cautious approach was to prevent the hole from back-filling with sand, and this was confirmed when logging tools passed safely to the base of the PQ string. After successfully completing the VSP run at 0805 h, the resistivity and acoustic sondes were run in the open-hole interval directly below the PQ string. Various attempts were then made to lift the PQ string, which was found to be stuck fast and likely to be held higher up in the hole. Attempts were made to cut the PQ pipe at 387 and 351 m DSF using a cutting tool, which finally succeeded at 286 m DSF before the HQ string and cutter also became stuck.

The HQ rods were worked through the night into 19 June in an effort to free the string. Meanwhile, a spare cutter was ordered from the shore. Other methods to free the HQ string were explored: a heavy tool was manufactured that was intended to break the tool joint connecting the cutter to the string, and a pinch bar was attached to the wireline and repeatedly hammered, through freefall, onto the top of the cutter. Neither method worked, although an increase in water flow, partial rotation, and slow pull-up of the HQ pipe was achieved. After further working of the rods, the HQ pipe was finally released and tripped with the cutter and box end missing.

While waiting for the spare cutter to arrive, a Bowen spear was deployed to try and pull the PQ. This was unsuccessful. Throughout the rest of the day, various attempts were made to free the PQ. Backing off was also attempted, but only resulted in disconnecting close to the drill floor and not below the seabed. The PQ rods continued to be worked overnight into 20 June. At 0600 h, the spare cutter arrived and shortly after 0900 h the PQ had been successfully cut just inside the casing and the HQ string and cutter were returned to deck. By 0930 h, what was left of the PQ string was recovered onto deck.

To free the stuck casing, the 6 inch casing cutter was modified by removing the centralizing ball bearings so that it was able to fit inside the buttress casing. The casing was successfully cut, allowing the team to recover the casing and rig down in anticipation of sailing to the next hole.

Transit to Hole M0029A

Preparations for transit to Hole M0029A commenced at 1455 h on 20 June 2009. This included setting a buoy to mark the location of the casing to be recovered from the seafloor. The jack-down procedure began at 1815 h, and the L/B Kayd moved off HOle M0028A at 1825 h, arriving at Hole M0029A at 1915 h. The legs were lowered to the seafloor, and preloading of the platform began at 1945 h. The drilling floor was opened up to operations personnel at 0630 h on 21 June to begin set-up of generators, powering up the containers, and preparation of the drilling floor.

Hole M0029A (proposed Site MAT-3)

During the morning of 21 June 2009, the ESO and drilling teams prepared the L/B Kayd for coring operations in Hole M0029A, and by 1030 h the casing was ready to be run in. The first three core runs had no recovery, and it was suspected that the soft sediment could not break through the microsphere bags. The microsphere bags were changed over to softer sample bags, and this, combined with a slight change in lithology, resulted in core collection in run four. For the rest of the day, continuous coring was completed to 19 m DSF, beyond which a switch was made to conducting a 3 m core run every 9 m. Coring continued to progress well for the rest of the day and through the night into 22 June.

After reaching 55 m DSF by 0530 h on 22 June, preparations were made to set the casing deeper into the seabed. The PQ string was pulled and the casing carefully rotated into the ground. To speed the process up, the L/B Kayd's jacking capability was used to assist in setting the casing. Rapid progress was made until the top drive sprung an oil leak and had to be stripped down. After setting had resumed, the casing was successfully positioned within a clay layer at ~11 m DSF. After reaming back down the hole, coring resumed by midnight and continued into Tuesday, 23 June, alternating with periods of reaming and circulating because of caving sands, down to 67.31 m DSF by 0850 h.

At 1155 h on 3 June, operations were halted when the PQ string became stuck and a twist-off of the core bit and reamer occurred. After tripping the PQ string, and fishing the bit and reaming shell back onto deck, a modified PQ outer core barrel and ALN core barrel was made up. The PQ pipe was run back in at 2230 h.

Early on 24 June, the base of the hole (69.36 m DSF) was reached and coring, with some open-holing, resumed at 76.47 m DSF. However, at 1150 h, sands with mud polymer patches caused the inner and outer core barrels to stick together, resulting in difficulty in rotating, flushing, and lifting the pipe. For the next 3 h, various attempts were made to release the string. The L/B Kayd was jacked down, taking the casing with it, allowing access to the PQ joint. Once the PQ was disconnected, the HQ was run to the base of the hole. Seawater was pumped down the HQ string and at 0400 h on Thursday, 25 June, rotation was reestablished. Six rod stands were removed with difficulty and the hole circulated. The PQ string was reamed to the base of the hole and the hole advanced by open-hole drilling to 117 m DSF where two core runs were made. This was followed by further open-hole drilling to 148 m DSF to the seismic reflection thought to be m1.

Open-hole drilling and spot coring continued throughout 26 June, apart from a short time when operations were suspended because of a lightning storm. By midnight the hole had advanced to 224 m DSF with 9 core runs. At 0100 h on 27 June, drilling operations were suspended for 1.5 h because of another lightning storm. After operations recommenced, open-hole drilling and spot coring continued until 1500 h, by which time the base of the hole was at 257 m DSF. The operation then switched to continuous coring.

Recovery was poor in loose sands, and various bit combinations were tried to improve recovery. Core recovery generally improved on 28 June, although it was still variable. At midnight, the base of the hole had advanced to 312 m DSF. Coring progressed steadily on 29 June, with recovery between 50% and 114%, reaching 352 m DSF by midnight. By midnight on 30 June, the base of the hole had advanced to 395 m DSF.

On 1 July, coring operations were suspended several times because of lightning storms. However, coring then continued from 395 m DSF with excellent core recovery on a steady basis, reaching the target depth of 756.65 m DSF by 1820 h on 11 July. On completion of coring, the rig floor was prepared for the logging program and the hole conditioned.

During the day, the VSP equipment, including the air guns, were prepared and tested. Marine mammal observations began 30 min prior to firing of the air guns, which started at 2050 h. The logging program began with through-pipe VSP from the base of the hole. Just after midnight a storm approached the platform and the lightning safety procedure was implemented. Through-pipe VSP operations were suspended at 0145 h (at ~200 m DSF). The deck was evacuated, and all personnel remained inside the accommodation module until it was safe to return to the deck at ~0400 h. Because of darkness at that time, it was not possible to restart the air gun, as the VSP permit stipulated start-up during daylight only. It was decided to abandon the through-pipe VSP operation and switch to through-pipe natural gamma wireline logging from 755 m DSF. The logging operation of the entire hole began at 0510 h and was completed successfully by 1215 h, although the winch motor was running very hot.

The next stage was to pull the pipe back to 600 m DSF and log the bottom, open-hole section of the hole. However, the drill pipe was stuck and required significant effort to free it. Alternating pulling pipe with reaming and flushing continued for most of the day without significant improvement in the condition of the hole. As by this time the pipe had been pulled back above the 600 m DSF level, the decision was made to continue the hole conditioning operation while pulling up to the next logging step at 480 m DSF, and from there try to log the bottom two sections of the hole. This depth was reached at 2245 h on 12 July.

Logging with the resistivity sonde began at 2350 h and was completed at 0300 h on 13 July, despite an intermittent electro/mechanical fault on the winch. The hole was conditioned, and the sonic sonde was lowered. At 624 m DSF, as the sonde was logging, the power supply to the winch tripped several times and eventually both the winch motor and control box burned out. The sonde was manually recovered to deck by pulling 600 m of wire up the drill pipe and was back on deck at 0905 h. The winch was removed from the drill floor and replaced with the back-up logging winch that had been used for VSP operations. Logging recommenced at 1110 h, first with the sonic and then with the magnetic susceptibility sonde. Between each logging run the drill hole was conditioned by circulating mud and rotating the pipes.

At 1925 h, the acoustic imaging sonde was connected to the winch cable on deck. No signal was detected, indicating a faulty tool, cable, or connector. Fault-finding efforts continued until 0200 h on 14 July with several faults being repaired and the tool replaced. On deployment of the tool at 0100 h, a further fault developed with the winch. Then, fatally, all contact was lost with the tool at 0235 h. The generator was also causing concern but kept functioning. The tool was retrieved by 0515 h, and fault finding restarted. An earth-leakage fault in the winch cable at the tool-connector end was finally identified. A section of cable was removed, and the end of the cable was reterminated. Once fixed and tested, the log run was successful, finishing at 1800 h.

Before tripping pipe up to the top of the next logging interval, the opportunity was taken to complete the through-pipe VSP that had been aborted on 12 July. Marine mammal observations started at 1700 h, and the top section of the hole (0–200 m DSF) was logged successfully, finishing at 2100 h. Once complete, the drill pipe was tripped back to the top of the next logging interval at 335 m DSF. It was noted that the drill pipes were very tight for the first five double stands.

On 14 July, the supply boat Sorensen Miller conducted a postdrill survey of proposed Site MAT-2D with side-scan sonar and magnetometer, prior to recovery of the casing at a later date.

On 15 July, tripping of the pipe was completed by 0150 h. The natural gamma sonde was lowered down the hole but was unable to progress very far. It punched through an upper bridge but could not penetrate deeper than 350 m DSF because of a second bridge. Logging was therefore conducted between 334 and 350 m DSF. A reaming operation started at 0825 h and was completed at 1845 h, having encountered a third bridge at 495 m DSF. Following this, the resistivity sonde was lowered down the hole at 1900 h. It was unable to penetrate the bridge at 350 m DSF despite reaming. Further reaming and probing indicated open-hole conditions below 357 m DSF. The decision was taken to ream down beneath both bridges and log the lower part of the scheduled logging interval. However, on reaming to this depth the probe revealed that a new bridge had formed at 400 m DSF. Further reaming continued to remove this bridge.

Logging with the resistivity, sonic, and magnetic susceptibility sondes continued until 1205 h on Thursday, 16 July. However, after connecting the acoustic imager sonde to the winch, an electrical problem in either the winch cable or connectors halted operations. Despite tripping pipe back to 352 m DSF to enable time for the repairs to be made, it became obvious that it would not be possible to complete repairs and finish the logging schedule within the remaining time because of transit time and the need to enter Atlantic City on a high tide. Therefore at 1425 h on 16 July, the decision was made to stop operations, trip all drill pipe and casing, and prepare the containers for transit back to Atlantic City.

Transit to Atlantic City, New Jersey

The L/B Kayd began preparations to depart from Hole M0029A at 0100 h on 17 July 2009, involving removing all power to the ESO containers, securing the drill floor and equipment, and moving all personnel into the accommodation block. Jacking-down of the platform began at 0200 h, and transit to Atlantic City commenced at 0315 h, with the vessel arriving alongside the Coast Guard Station quayside at 1545 h.

Demobilization of the ESO containers and equipment continued on 17 and 18 July, with the containers being lifted onto the quayside on 18 July. On 19 July, all containers were collected for shipping back to Europe, with the wireline tools being collected on Monday, 20 July. All ESO staff departed on 20 July.

Onshore Science Party, Bremen

The cores and samples collected offshore New Jersey were transported under refrigeration to the IODP Bremen Core Repository and Laboratories in the MARUM building on the campus of Bremen University (Germany). Frozen microbiology samples were forwarded from Bremen to scientists' laboratories in Germany and France by special courier service. NGR logging and thermal conductivity measurements were acquired before the start of the OSP. Further analytical laboratories were accessed by agreement with the Department of Geosciences (geochemistry, paleomagnetics, mineralogy [X-ray diffraction], and hydrofluoric acid laboratories) and the Center for Marine Environmental Sciences (MARUM) (physical properties, micropaleontology, and nondestructive core logging laboratories) at Bremen University.

During the Expedition 313 Onshore Science Party (6 November–4 December 2009), the cores were described in detail and minimum and some standard measurements were made (Table T2). In addition, sampling for postcruise scientific research was also undertaken.